4–76. The number of newspapers demanded daily in a large metropolitan area is
believed to be an approximately normally distributed random variable. If more
newspapers are demanded than are printed, the paper suffers an opportunity loss,in that it could have sold more papers, and a loss of public goodwill. On the other
hand, if more papers are printed than will be demanded, the unsold papers are
returned to the newspaper office at a loss. Suppose that management believes that
guarding against the first type of error, unmet demand, is most important and
would like to set the number of papers printed at a level such that 75% of the time,
demand for newspapers will be lower than that point. How many papers should be
printed daily if the average demand is 34,750 papers and the standard deviation of
demand is 3,560?

Answer by stanbon(75887)

Suppose that management believes that
guarding against the first type of error, unmet demand, is most important and
would like to set the number of papers printed at a level such that 75% of the time, demand for newspapers will be lower than that point. How many papers should be printed daily if the average demand is 34,750 papers and the standard deviation of demand is 3,560?
-----
Find the z-value that has a left tail of 0.75.
invNorm(0.75) = 0.6745
---
Find the corresponding x-value.
----
x = zs+u
x = 0.6745*3560+34750 = 37151 papers printed daily
==================
Cheers,
Stan H.